# Uncovering the Mechanism of Chiral Three-Nucleon Force in Driving Spin-Orbit Splitting

### v1

@@ Abstract
The three-nucleon force (3NF) is crucial in shaping the shell structure of atomic nuclei, particularly impacting the enhancement of spin-orbit (SO) splitting, especially in nuclei with significant deviations from stability. Despite its importance, the specific mechanisms driving this enhancement remain unclear. In this study, we introduce a decomposition scheme based on the rank of irreducible tensors forming the 3NF, derived from chiral effective field theory at next-to-next-to-leading order, to elucidate their influence on SO splitting. Within the shell-model framework, our analysis reveals that the rank-1 component of the 3NF is the primary factor enlarging the energy gap between the $0p_{3/2}$ and $0p_{1/2}$ single-particle levels in $p$-shell nuclei, while the rank-2 component makes a subdominant contribution. We also remark on the antisymmetry of the rank-1 3NF, as a means to clarify the disentanglement of quantum spin states. This study lays the groundwork for further exploration into this field toward a microscopic understanding of the 3NF impact on the nuclear shell structure.
Introduction.— The understanding of the structure and dynamics of atomic nuclei hinges not only on the intricate nature of two-nucleon forces (2NFs) but also on the significant impact of three-nucleon forces (3NFs) (see for instance Ref. [1]). Since Primakoff and Holstein [2] argued for the importance of many-body forces in quantum systems, considerable efforts have been dedicated to studying 3NFs in nuclear physics. Initiating with the Fujita–Miyazawa force [3], various methods have been proposed to construct 3NFs based on pion exchanges (see Refs. [4, 5] for the comparison of 3NFs). Recent advancements rooted in studies by Weinberg [6] and van Kolck with coauthors [7, 8] on 3NFs from a chiral Lagrangian have notably progressed with the advent of chiral effective field theory (EFT) [9, 10, 11]. This theory offers a systematic and unified approach to constructing 2NFs and 3NFs on equal footing [6, 7, 11].
@@
Method.— The chiral 3NF at N2LO contains three terms; the two-pion ($2\pi$) exchange term, one-pion ($1\pi$) exchange plus contact term, and contact (ct) term, expressed by the left, middle, and right diagrams of Fig. 1, respectively. Furthermore, the $2\pi$ term comprises three distinct constituents associated with the three low-energy constants (LECs), $c_{1}$, $c_{3}$, and $c_{4}$, which initially appear at the 2NF level and are determined using the data of the two-nucleon systems. The potential of each term of the 3NF can be schematically written as
$v_{3N}^{(\alpha)}$ $=\sum_{i\neq j\neq k}v^{(\alpha)}\!\left(\bm{\tau}_{i},\bm{\tau}_{j},\bm{\tau}_{k}\right)$
×$\times w^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right),$ (1)
where $v^{(\alpha)}$ is the isospin part and $w^{(\alpha)}$ represents the spin-momentum dependent part. The letter $\alpha$ stands for ct, $1\pi$, or one of the three constituents of the $2\pi$ term, and the indices $i,j,k\in\{1,2,3\}$ specify the nucleon. The spin and isospin operators of the nucleon ${i}$ are denoted by the Pauli matrices, $\bm{\sigma}_{i}$ and $\bm{\tau}_{i}$, respectively. The momentum transfer is given by ${\bm{q}_{i}=\bm{p}_{i}^{\prime}-\bm{p}}$, where $\bm{p}_{i}$ ($\bm{p}_{i}^{\prime}$) is the initial (final) momentum of the nucleon $i$.
@@
$w^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right)$
$\quad=w_{\mathrm{pro}}^{(\alpha)}(q_{i},q_{j})\sum_{\lambda}\mathcal{O}_{\lambda}^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right),$ (2)
$\mathcal{O}_{\lambda}^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$
$\quad=A_{\lambda}\left[\mathcal{M}_{\lambda}^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k}\right)\otimes\mathcal{N}_{\lambda}^{(\alpha)}\!\left(\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)\right]_{00}.$ (3)
The irreducible tensors of rank $\lambda$ of the spin and momentum parts are expressed by $\mathcal{M}_{\lambda}^{(\alpha)}$ and $\mathcal{N}_{\lambda}^{(\alpha)}$, respectively, with the coefficient $A_{\lambda}$ and ${\hat{\bm{q}}_{i}=\bm{q}_{i}/q_{i}}$. Equation (3) denotes the product of the two tensors forming the scalar ${[\cdots\otimes\cdots]_{00}}$. The explicit expressions of all quantities appearing in Eqs. (2) and (3), including the function $w_{\mathrm{pro}}^{(\alpha)}$, are presented in the Supplemental Material [48], along with the standard expressions of the three terms of the chiral 3NF at N2LO.
@@
Moreover, referring to Ref. [43], the effective shell-model Hamiltonian for the $0p$-shell valence space outside the doubly magic ${}^{4}\mathrm{He}$ was derived within the framework of the many-body perturbation theory [50, 51, 52], accounting for one- and two-body diagrams up to third order for two-body interaction, as well as one- and two-body diagrams up to first order for the 3NF. These 3NF diagrams correspond to the normal-ordered one-body (NO1B) and normal-ordered two-body (NO2B) terms of the usual expression of the 3NF in the normal-ordered form [53]. However, in contrast to Ref. [43], we did not include folded diagrams for the 3NF [27]. In this study, we added the NO1B and NO2B terms of the 3NF, respectively, to the single-particle energies (SPEs) and the two-body matrix elements of the effective shell-model Hamiltonian derived from the 2NF plus Coulomb force.
The omission of the 3NF-folded diagrams implies that changes in the SO splitting induced by the 3NF in systems with one-particle outside a doubly closed core can be directly related to the NO1B term, while the NO2B one contributes to the evolution of the SO splitting as a function of the number of valence nucleons. Nevertheless, we have verified that the quantities relevant to the SO splitting, which are discussed in this Letter, namely, the SPEs and effective SPEs (ESPEs) [54], are not significantly affected by this choice. See the Supplemental Material [48] for the definition of the ESPEs.
@@
The mechanism through which the rank-1 3NF enlarges the gap of the SPEs is determined by the characteristic of the $c_{3}$ term of the $2\pi$-exchange process. The $c_{3}$ potential with $\lambda=1$ depends on ${\left(\bm{\tau}_{i}\cdot\bm{\tau}_{j}\right)\left(\bm{\sigma}_{i}\times\bm{\sigma}_{j}\right)\cdot\left(\hat{\bm{q}}_{i}\times\hat{\bm{q}}_{j}\right)\left(\hat{\bm{q}}_{i}\cdot\hat{\bm{q}}_{j}\right)}$, which is given by Eq. (29), and generates the one-body SO potential, as expressed by Eq. (4.23) of Ref. [40]. In particular, this explanation also applies to the SO splitting observed in the ESPEs reported herein.
@@
Subsequently, using the SPEs reported above, we consider the evolution of the $0p$-SO splitting. With this perspective, we investigate the ESPEs, which represent the single-particle states generated by the mean field incorporating the effects from other nucleons outer the inert core. The occupation numbers required in the calculation of the ESPEs are reported in Table 2 [48]; these values are obtained in the shell-model framework using the kshell code [57].
@@
Conclusion and perspectives.— We introduced the decomposition of the chiral 3NF at N2LO in terms of the rank of irreducible tensors. Based on this technique, we demonstrated the predominant influence of the rank-1 3NF on the SO splitting in the $p$-shell nuclei. This study offers a fresh perspective on the mechanisms shaping shell structure driven by nuclear forces.
The rank-1 3NF is potentially accessible through specific observables. As reported in Refs. [61, 62], linear combinations of the vector analyzing powers in proton–deuteron scattering are particularly sensitive to the rank-1 3NF. Consequently, these observables have the potential to constrain the chiral 3NF, thereby facilitating quantitative conclusions regarding its contributions to the SO splitting in the future.
@@
In conclusion, we remark on the antisymmetry of the rank-1 3NF. The antisymmetric nature manifests in the two-body subsystems of the complete three-body system, owing to the operator $\bm{\sigma}_{i}\times\bm{\sigma}_{j}$ of the $c_{1}$ and $c_{3}$ terms [see Eqs. (25) and (29)]. This operator induces singlet–triplet mixing of the spin states in the two-body subsystems, akin to the spin canting of magnetic ions driven by the Dzyaloshinsky–Moriya interaction [63, 64, 65]. Consequently, the rank-1 3NF disentangles the quantum entanglement manifested as spin-singlet states, which are expressed as the superposition of the product states of the spin-up and spin-down components. While quantum entanglement in nuclei has been intensively studied through nucleon–nucleon correlations [66, 67, 68, 69, 70, 71, 72, 73, 74], exploring the disentanglement prompted by the 3NF presents an intriguing avenue for future research.
Acknowledgements.
@@
The authors are deeply grateful to L. Coraggio, N. Itaco, Y. Kanada-En’yo, M. Kimura, K. Nomura, Y. Taniguchi, Y. Tanimura, and Y. Yamaguchi for fruitful discussions and useful comments. This work was supported by JSPS KAKENHI Grant Number JP21K13919 and JST ERATO Grant Number JPMJER2304. The calculations were carried out using the computer facilities at Yukawa Institute for Theoretical Physics, Kyoto University.
Appendix A Supplemental Online Material for:
Uncovering Mechanism of Chiral Three-Nucleon Force in Driving Spin-Orbit Splitting
Tokuro Fukui1,2, Giovanni De Gregorio3,4, and Angela Gargano4
1Faculty of Arts and Science, Kyushu University, Fukuoka 819-0395, Japan
2RIKEN Nishina Center, Wako 351-0198, Japan
3Dipartimento di Matematica e Fisica, Università degli Studi della Campania
“Luigi Vanvitelli”, viale Abramo Lincoln 5-I-81100 Caserta, Italy
4Istituto Nazionale di Fisica Nucleare, Complesso Universitario di Monte
S. Angelo, Via Cintia, I-80126 Napoli, Italy
In this Supplemental Material, we present the explicit forms of the potential of the three-nucleon force decomposed by the irreducible tensors. We also give the definition of the effective single-particle energy to clearly show how the three-nucleon force enter there. The occupation numbers relevant to the calculations of the effective single-particle energies are also reported.
@@ Appendix A Supplemental Online Material for:
Appendix B Irreducible-tensor decomposition of chiral three-nucleon force
B.0.1 Standard expression of potential
@@
Here we explicitly show the irreducible-tensor decomposition of the potential of the three-nucleon force (3NF) derived within the chiral effective field theory at at next-to-next-to-leading order (N2LO).
@@
$v_{3N}^{\rm(ct)}$ $=\frac{c_{E}}{2f_{\pi}^{4}\Lambda_{\chi}}\sum_{i\neq j\neq k}\bm{\tau}_{i}\cdot\bm{\tau}_{j},$ (4)
$v_{3N}^{(1\pi)}$ $=-\frac{g_{A}c_{D}}{8f_{\pi}^{4}\Lambda_{\chi}}\sum_{i\neq j\neq k}\bm{\tau}_{i}\cdot\bm{\tau}_{j}\frac{\left(\bm{\sigma}_{j}\cdot\bm{q}_{j}\right)\left(\bm{\sigma}_{i}\cdot\bm{q}_{j}\right)}{q_{j}^{2}+m_{\pi}^{2}},$ (5)
$v_{3N}^{(c_{1})}$ $=-\frac{g_{A}^{2}c_{1}m_{\pi}^{2}}{2f_{\pi}^{4}}\sum_{i\neq j\neq k}\bm{\tau}_{i}\cdot\bm{\tau}_{j}$
×$\times\frac{\left(\bm{\sigma}_{i}\cdot\bm{q}_{i}\right)\left(\bm{\sigma}_{j}\cdot\bm{q}_{j}\right)}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)},$ (6)
$v_{3N}^{(c_{3})}$ $=\frac{g_{A}^{2}c_{3}}{4f_{\pi}^{4}}\sum_{i\neq j\neq k}\bm{\tau}_{i}\cdot\bm{\tau}_{j}$
×$\times\frac{\left(\bm{\sigma}_{i}\cdot\bm{q}_{i}\right)\left(\bm{\sigma}_{j}\cdot\bm{q}_{j}\right)}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}\bm{q}_{i}\cdot\bm{q}_{j},$ (7)
$v_{3N}^{(c_{4})}$ $=\frac{g_{A}^{2}c_{4}}{8f_{\pi}^{4}}\sum_{i\neq j\neq k}\left(\bm{\tau}_{i}\times\bm{\tau}_{j}\right)\cdot\bm{\tau}_{k}$
×$\times\frac{\left(\bm{\sigma}_{i}\cdot\bm{q}_{i}\right)\left(\bm{\sigma}_{j}\cdot\bm{q}_{j}\right)}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}\left(\bm{q}_{i}\times\bm{q}_{j}\right)\cdot\bm{\sigma}_{k},$ (8)
where the nucleon spin (isospin) operator is expressed by the Pauli matrices $\bm{\sigma}_{i}$ ($\bm{\tau}_{i}$). The momentum transfer is defined by ${\bm{q}_{i}=\bm{p}_{i}^{\prime}-\bm{p}}$, with $\bm{p}_{i}$ and $\bm{p}_{i}^{\prime}$ the initial and final momenta of the nucleon $i$, respectively. The LECs $c_{E}$ and $c_{D}$ are associated with the contact and $1\pi$ terms, respectively. The above potentials contain also the pion-decay constant $f_{\pi}$, the chiral symmetry breaking scale $\Lambda_{\chi}$, the axial vector coupling constant $g_{A}$, and the average pion mass $m_{\pi}$.
@@
$v_{3N}^{(\alpha)}$ $=\sum_{i\neq j\neq k}v^{(\alpha)}\!\left(\bm{\tau}_{i},\bm{\tau}_{j},\bm{\tau}_{k}\right)$
×$\times w^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right).$ (9)
@@
$v^{(\mathrm{ct})}\!\left(\bm{\tau}_{i},\bm{\tau}_{j},\bm{\tau}_{k}\right)$ $=v^{(1\pi)}\!\left(\bm{\tau}_{i},\bm{\tau}_{j},\bm{\tau}_{k}\right)$
$=v^{(c_{1})}\!\left(\bm{\tau}_{i},\bm{\tau}_{j},\bm{\tau}_{k}\right)$
$=v^{(c_{3})}\!\left(\bm{\tau}_{i},\bm{\tau}_{j},\bm{\tau}_{k}\right)$
$=\bm{\tau}_{i}\cdot\bm{\tau}_{j},$ (10)
$v^{(c_{4})}\!\left(\bm{\tau}_{i},\bm{\tau}_{j},\bm{\tau}_{k}\right)$ $=\left(\bm{\tau}_{i}\times\bm{\tau}_{j}\right)\cdot\bm{\tau}_{k}.$ (11)
B.0.2 General expression of irreducible-tensor decomposition

### v2

@@ Abstract
The three-nucleon force (3NF) is crucial in shaping the shell structure of atomic nuclei, particularly impacting the enhancement of spin-orbit (SO) splitting, especially in nuclei with significant deviations from stability. Despite its importance, the specific mechanisms driving this enhancement remain unclear. In this study, we introduce a decomposition scheme based on the rank of irreducible tensors forming the 3NF, derived from chiral effective field theory at next-to-next-to-leading order, to elucidate their influence on SO splitting. Within the shell-model framework, our analysis reveals that the rank-1 component of the 3NF is the primary factor enlarging the energy gap between the $0p_{3/2}$ and $0p_{1/2}$ single-particle levels in $p$-shell nuclei, while the rank-2 component makes a subdominant contribution. Since the rank-1 component originates exclusively from the $2\pi$-exchange 3NF, our finding will not depend on the choice of the low-energy constants of contact terms. We also remark on the antisymmetry of the rank-1 3NF, which can affect the quantum entanglement of spin states. This study lays the groundwork for further exploration into this field toward a microscopic understanding of the 3NF impact on the nuclear shell structure.
keywords:
Three-nucleon force , Spin-orbit splitting , Chiral effective field theory , Shell model
journal: Physics Letters B
1 Introduction
The understanding of the structure and dynamics of atomic nuclei hinges not only on the intricate nature of two-nucleon forces (2NFs) but also on the significant impact of three-nucleon forces (3NFs) (see for instance Ref. [1]). Since Primakoff and Holstein [2] argued for the importance of many-body forces in quantum systems, considerable efforts have been dedicated to studying 3NFs in nuclear physics. Initiating with the Fujita–Miyazawa force [3], various methods have been proposed to construct 3NFs based on pion exchanges (see Refs. [4, 5] for the comparison of 3NFs). Recent advancements rooted in studies by Weinberg [6] and van Kolck with coauthors [7, 8] on 3NFs from a chiral Lagrangian have notably progressed with the advent of chiral effective field theory (EFT) [9, 10, 11]. This theory offers a systematic and unified approach to constructing 2NFs and 3NFs on equal footing [6, 7, 11].
@@ 1 Introduction
2 Method
The chiral 3NF at N2LO contains three terms; the two-pion ($2\pi$) exchange term, one-pion ($1\pi$) exchange plus contact term, and contact (ct) term, expressed by the left, middle, and right diagrams of Fig. 1, respectively. Furthermore, the $2\pi$ term comprises three distinct constituents associated with the pion-nucleon low-energy constants (LECs), $c_{1}$, $c_{3}$, and $c_{4}$, which also appear at the 2NF level. The potential of each term of the 3NF can be schematically written as
$v_{3N}^{(\alpha)}$ $=\sum_{i\neq j\neq k}v^{(\alpha)}\!\left(\boldsymbol{\tau}_{i},\boldsymbol{\tau}_{j},\boldsymbol{\tau}_{k}\right)w^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right),$ (1)
where $v^{(\alpha)}$ is the isospin part and $w^{(\alpha)}$ represents the spin-momentum dependent part. The letter $\alpha$ stands for ct, $1\pi$, or one of the three constituents of the $2\pi$ term, and the indices $i,j,k\in\{1,2,3\}$ specify the nucleon. The spin and isospin operators of the nucleon ${i}$ are denoted by the Pauli matrices, $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\tau}_{i}$, respectively. The momentum transfer is given by ${\boldsymbol{q}_{i}=\boldsymbol{p}_{i}^{\prime}-\boldsymbol{p}}$, where $\boldsymbol{p}_{i}$ ($\boldsymbol{p}_{i}^{\prime}$) is the initial (final) momentum of the nucleon $i$.
@@ 2 Method
$w^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right)$
$\quad=w_{\mathrm{pro}}^{(\alpha)}(q_{i},q_{j})\sum_{\lambda}\mathcal{O}_{\lambda}^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right),$ (2)
$\mathcal{O}_{\lambda}^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
$\quad=A_{\lambda}\left[\mathcal{M}_{\lambda}^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k}\right)\otimes\mathcal{N}_{\lambda}^{(\alpha)}\!\left(\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)\right]_{00}.$ (3)
The irreducible tensors of rank $\lambda$ of the spin and momentum parts are expressed by $\mathcal{M}_{\lambda}^{(\alpha)}$ and $\mathcal{N}_{\lambda}^{(\alpha)}$, respectively, with the coefficient $A_{\lambda}$ and ${\hat{\boldsymbol{q}}_{i}=\boldsymbol{q}_{i}/q_{i}}$. Equation (3) denotes the product of the two tensors forming the scalar ${[\cdots\otimes\cdots]_{00}}$. The explicit expressions of all quantities appearing in Eqs. (2) and (3), including the function $w_{\mathrm{pro}}^{(\alpha)}$, are presented in A, along with the standard expressions of the three terms of the chiral 3NF at N2LO. Equation (2) is based on the decomposition of the three-body potentials rather than the corresponding matrix elements. In contrast, the spin-tensor decomposition [48, 49, 50] of two-body matrix elements (2BMEs) is widely used for analyzing tensorial structure of 2NFs.
@@
Moreover, referring to Ref. [43], the effective shell-model Hamiltonian for the $0p$-shell valence space outside the doubly magic ${}^{4}\mathrm{He}$ was derived within the framework of the many-body perturbation theory [52, 53, 54]. Specifically, the effective shell-model Hamiltonian is expressed by way of the Kuo-Lee-Ratcliff folded-diagram expansion [55] in terms of the $\hat{Q}$-box vertex function, which was constructed by including one- and two-body diagrams up to third order for two-body interaction, as well as one- and two-body diagrams up to first order for the 3NF. These 3NF diagrams correspond to the normal-ordered one-body (NO1B) and normal-ordered two-body (NO2B) terms of the usual expression of the 3NF in the normal-ordered form [56]. This so-called normal-ordering approximation helps manage the complexity of fully treating the 3NF, and has been validated in light- and medium-mass nuclei [56, 24, 57]. Within the shell-model framework, the validity of this approximation has been tested for $0p$-shell nuclei through benchmark calculations comparing shell-model results [43] with those from the ab initio no-core shell model [58, 12]. In contrast to Ref. [43], in the present calculations we did not include folded diagrams for the 3NF [27], i.e., we added the NO1B and NO2B terms of the 3NF, respectively, to the single-particle energies (SPEs) and the 2BMEs of the effective shell-model Hamiltonian derived from the 2NF plus Coulomb force.
The omission of the 3NF-folded diagrams implies that changes in the SO splitting induced by the 3NF in systems with one-particle outside a doubly closed core can be directly related to the NO1B term, while the NO2B one contributes to the evolution of the SO splitting as a function of the number of valence nucleons. In doing so, we neglect some folding contributions related to the 3NF. However, we have verified that such contributions are not significant and do not affect the quantities relevant to the SO splitting, which are discussed in this Letter, namely, the SPEs and effective SPEs (ESPEs) [59]. See B for the definition of the ESPEs. The little relevance of the 3NF in the folding expansion, expressed as products of $\hat{Q}$-box and its derivatives [53], is essentially due to the zero value of the first-order-diagram derivatives, which do not depend in fact on the energy.
3 The finding: Importance of rank-1 3NF
@@
The mechanism through which the rank-1 3NF enlarges the gap of the SPEs is determined by the characteristic of the $c_{3}$ term of the $2\pi$-exchange process. The $c_{3}$ potential with $\lambda=1$ depends on ${\left(\boldsymbol{\tau}_{i}\cdot\boldsymbol{\tau}_{j}\right)\left(\boldsymbol{\sigma}_{i}\times\boldsymbol{\sigma}_{j}\right)\cdot\left(\hat{\boldsymbol{q}}_{i}\times\hat{\boldsymbol{q}}_{j}\right)\left(\hat{\boldsymbol{q}}_{i}\cdot\hat{\boldsymbol{q}}_{j}\right)}$, which is given by Eq. (29), and generates the one-body SO potential, as expressed by Eq. (4.23) of Ref. [40]. In particular, this explanation also applies to the SO splitting observed in the ESPEs reported herein.
@@
Subsequently, using the SPEs reported above, we consider the evolution of the $0p$-SO splitting. With this perspective, we investigate the ESPEs, which represent the single-particle states generated by the mean field incorporating the effects from other nucleons outer the inert core. In fact, changes in the ESPEs induced by the effects of the 3NF on the monopole component of the interaction are crucial to describe the observed spectroscopic properties of nuclei (see Ref. [27] and references therein). The occupation numbers required in the calculation of the ESPEs are reported in Table 2 in B; these values are obtained in the shell-model framework using the kshell code [62].
@@
4 Conclusion and perspectives
We introduced the decomposition of the chiral 3NF at N2LO in terms of the rank of irreducible tensors. Based on this technique, we demonstrated the predominant influence of the rank-1 3NF on the SO splitting in the $p$-shell nuclei. This conclusion will not be affected by the choice of the contact LECs since the rank-1 component originates exclusively from the $2\pi$-exchange 3NF. In this connection, it would be valuable to investigate whether the rank-1 3NF we use can account for specific observables of proton–deuteron scattering, namely, linear combinations of the vector analyzing powers, which are particularly sensitive to the rank-1 3NF as reported in Refs. [67, 68].
This study offers a fresh perspective on the mechanisms shaping shell structure driven by nuclear forces. However, to assess the robustness of our conclusions, it is essential to extend our analysis to heavier mass regions. In a near future, we plan to investigate the contribution of the different rank components of the 3NF in $pf$-shell nuclei, for which the global effect of the 3NF was already evidenced [25].
@@ 4 Conclusion and perspectives
In conclusion, we remark on the antisymmetry of the rank-1 3NF. The antisymmetric nature manifests in the two-body subsystems of the complete three-body system, owing to the operator $\boldsymbol{\sigma}_{i}\times\boldsymbol{\sigma}_{j}$ of the $c_{1}$ and $c_{3}$ terms [see Eqs. (25) and (29)]. This operator induces singlet–triplet mixing of the spin states in the two-body subsystems, akin to the spin canting of magnetic ions driven by the Dzyaloshinsky–Moriya interaction [69, 70, 71]. Consequently, the rank-1 3NF can change the quantum entanglement manifested as the superposition of the product states of the spin-up and spin-down components. While quantum entanglement in nuclei has been intensively studied through nucleon–nucleon correlations [72, 73, 74, 75, 76, 77, 78, 79, 80], exploring the change in quantum entanglement prompted by the 3NF presents an intriguing avenue for future research.
Acknowledgements
@@
The authors are deeply grateful to L. Coraggio, N. Itaco, Y. Kanada-En’yo, M. Kimura, K. Nomura, Y. Taniguchi, Y. Tanimura, and Y. Yamaguchi for fruitful discussions and useful comments. This work was supported by JSPS KAKENHI Grant Numbers JP21K13919 and JP23KK0250, as well as JST ERATO Grant Number JPMJER2304. The calculations were carried out using the computer facilities at Yukawa Institute for Theoretical Physics, Kyoto University
@@
Appendix A Irreducible-tensor decomposition of chiral three-nucleon force
A.1 Standard expression of potential
@@
Here we explicitly show the irreducible-tensor decomposition of the potential of the three-nucleon force (3NF) derived within the chiral effective field theory at next-to-next-to-leading order (N2LO).
@@
$v_{3N}^{\rm(ct)}$ $=\frac{c_{E}}{2f_{\pi}^{4}\Lambda_{\chi}}\sum_{i\neq j\neq k}\boldsymbol{\tau}_{i}\cdot\boldsymbol{\tau}_{j},$ (4)
$v_{3N}^{(1\pi)}$ $=-\frac{g_{A}c_{D}}{8f_{\pi}^{4}\Lambda_{\chi}}\sum_{i\neq j\neq k}\boldsymbol{\tau}_{i}\cdot\boldsymbol{\tau}_{j}\frac{\left(\boldsymbol{\sigma}_{j}\cdot\boldsymbol{q}_{j}\right)\left(\boldsymbol{\sigma}_{i}\cdot\boldsymbol{q}_{j}\right)}{q_{j}^{2}+m_{\pi}^{2}},$ (5)
$v_{3N}^{(c_{1})}$ $=-\frac{g_{A}^{2}c_{1}m_{\pi}^{2}}{2f_{\pi}^{4}}\sum_{i\neq j\neq k}\boldsymbol{\tau}_{i}\cdot\boldsymbol{\tau}_{j}\frac{\left(\boldsymbol{\sigma}_{i}\cdot\boldsymbol{q}_{i}\right)\left(\boldsymbol{\sigma}_{j}\cdot\boldsymbol{q}_{j}\right)}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)},$ (6)
$v_{3N}^{(c_{3})}$ $=\frac{g_{A}^{2}c_{3}}{4f_{\pi}^{4}}\sum_{i\neq j\neq k}\boldsymbol{\tau}_{i}\cdot\boldsymbol{\tau}_{j}\frac{\left(\boldsymbol{\sigma}_{i}\cdot\boldsymbol{q}_{i}\right)\left(\boldsymbol{\sigma}_{j}\cdot\boldsymbol{q}_{j}\right)}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}\boldsymbol{q}_{i}\cdot\boldsymbol{q}_{j},$ (7)
$v_{3N}^{(c_{4})}$ $=\frac{g_{A}^{2}c_{4}}{8f_{\pi}^{4}}\sum_{i\neq j\neq k}\left(\boldsymbol{\tau}_{i}\times\boldsymbol{\tau}_{j}\right)\cdot\boldsymbol{\tau}_{k}$
×$\times\frac{\left(\boldsymbol{\sigma}_{i}\cdot\boldsymbol{q}_{i}\right)\left(\boldsymbol{\sigma}_{j}\cdot\boldsymbol{q}_{j}\right)}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}\left(\boldsymbol{q}_{i}\times\boldsymbol{q}_{j}\right)\cdot\boldsymbol{\sigma}_{k},$ (8)
where the nucleon spin (isospin) operator is expressed by the Pauli matrices $\boldsymbol{\sigma}_{i}$ ($\boldsymbol{\tau}_{i}$). The momentum transfer is defined by ${\boldsymbol{q}_{i}=\boldsymbol{p}_{i}^{\prime}-\boldsymbol{p}}$, with $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{i}^{\prime}$ the initial and final momenta of the nucleon $i$, respectively. The LECs $c_{E}$ and $c_{D}$ are associated with the contact and $1\pi$ terms, respectively. The above potentials contain also the pion-decay constant $f_{\pi}$, the chiral symmetry breaking scale $\Lambda_{\chi}$, the axial vector coupling constant $g_{A}$, and the average pion mass $m_{\pi}$.
@@
$v_{3N}^{(\alpha)}$ $=\sum_{i\neq j\neq k}v^{(\alpha)}\!\left(\boldsymbol{\tau}_{i},\boldsymbol{\tau}_{j},\boldsymbol{\tau}_{k}\right)w^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right).$ (9)
@@
$v^{(\mathrm{ct})}\!\left(\boldsymbol{\tau}_{i},\boldsymbol{\tau}_{j},\boldsymbol{\tau}_{k}\right)$ $=v^{(1\pi)}\!\left(\boldsymbol{\tau}_{i},\boldsymbol{\tau}_{j},\boldsymbol{\tau}_{k}\right)$
$=v^{(c_{1})}\!\left(\boldsymbol{\tau}_{i},\boldsymbol{\tau}_{j},\boldsymbol{\tau}_{k}\right)$
$=v^{(c_{3})}\!\left(\boldsymbol{\tau}_{i},\boldsymbol{\tau}_{j},\boldsymbol{\tau}_{k}\right)$
$=\boldsymbol{\tau}_{i}\cdot\boldsymbol{\tau}_{j},$ (10)
$v^{(c_{4})}\!\left(\boldsymbol{\tau}_{i},\boldsymbol{\tau}_{j},\boldsymbol{\tau}_{k}\right)$ $=\left(\boldsymbol{\tau}_{i}\times\boldsymbol{\tau}_{j}\right)\cdot\boldsymbol{\tau}_{k}.$ (11)
A.2 General expression of irreducible-tensor decomposition